OpenStudy (anonymous):

2 log 50= 3 log 25 +log (x-2)

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Use: \[a \log(m) = \log(m^a)\]

OpenStudy (anonymous):

And then you can use: \[\log(m) + \log(n) = \log(mn)\]

OpenStudy (anonymous):

Refer to the attachment.